(4) Consider a market with the following demand and supply curves:

                                                 Q (p) = 20 – 2P               

                                                 Q (p) = - 10 + 3P

(b)Suppose the government imposed a sales tax of $0.80 per unit of output sold, find the price paid by the consumer, the price received by the supplier, the equilibrium quantity transacted, and the total tax revenue received by the government.


(c)What do you think is the purpose for such tax policy? In your explanation, include a brief discussion of the importance of price elasticity of demand in the choice of commodities to be taxed in order to achieve specific policy goal(s).


1
Expert's answer
2021-12-14T11:21:28-0500

Part-B)

  • The government imposed a sales tax of $0.80 per unit of output sold, find the price paid by the consumer, the price received by the supplier, the equilibrium quantity transacted, and the total tax revenue received by the government:

Now, when the sales tax is imposed, the supply equation will be as:

"Qs = −10 + 3P\\

Qs = −10 + 3(P − 0.80)"

Thus, the new equilibrium structure will be

"QD = QS\\20 − 2P = −10 + 3(P − 0.80)\\20 − 2p = −10 + 3P − 2.4\\20 + 10 − 2P = 3P − 2.4\\30 − 2P = 3P − 2.4\\30 + 2.4 = 2P + 3P\\32.4 = 5P\\P =\frac{ 32.4}{5}\\P = \$6.48 \\

Now, \\

Qd = 20 − 2P\\Qd = 20 − 2(6.48)\\Qd = 20 − 12.96\\Qd = 7.04"


A) Here the price paid by the consumer is $6.48

Now, 

Price Received = $6.48 − $0.80

Price Received = $5.68

Price Received = $6.48 - $0.80Price Received = $5.68

B) Price received by the supplier is $5.68

C) Equilibrium quantity is 7.04 units

And, 

Tax Collected = $0.80 × 7.04

Tax Collected = $5.632

D) The total tax revenue received by the government is $5.632

Part-C)

  • What do you think is the purpose for such a tax policy:

Answer:

Now, the main aim towards collecting or generating the tax policy is firstly towards the betterment of society growth (development process like households, medicals, schools, etc.) that increases the growth level.

Under the price elasticity of demand, the tax policy will be beneficial when the demand is inelastic i.e. an increase in price leads to the same demand i.e. demand remains unchanged. Therefore, the government can and may implement the taxes towards higher revenue and can increase the spending level for healthy GDP.